A transformer has 500. primary turns and 10.0 secondary turns. (a) If Vp is 120.V(rms), what is Vs? If the secondary now has a resistive load of 10.0Ω, what is the current in the (b) secondary and (c) primary?
Answer #1

a) for the secondary voltage

Vs/Vp = Ns/Np

Vs/120 = 10/500

Vs = 2.4 V

the secondary voltage Vs is 2.4 V

b)

current in secondary = Vs/Resistance

current in secondary = 2.4/(10)

current in secondary = 0.24 A

the current in secondary is 0.24 A

c)

for the priamry current

as Vs * Is = Vp * Ip

0.24 * 2.4 = 120 * Ip

Ip = 0.0048 A

the current in the primary is 0.0048 A